Using graphical method, solve the game whose pay-off matrix is given as:

Player B

I II III IV

Player A I 1 3 -3 7

II 2 5 4 -6




1
Expert's answer
2021-06-15T13:45:08-0400

Column MiniMax =2

Row MaxiMin =-3


1.

The point of optimal solution occurs at the intersection of two lines:

"E_1=p_1-3p_2"

"E_3=2p_1+4p_2"

"p_1-3p_2=2p_1+4p_2"

"p_2=1-p_1"

"1-4p_2=2+2p_2"

"p_2=-1\/6" ",p_1=7\/6"

"V=7\/6+3\/6=5\/3"


2.

The point of optimal solution occurs at the intersection of two lines:

"L_1=q_1+2q_2"

"L_3=-3q_1+4q_2"

"q_1+2q_2=-3q_1+4q_2"

"q_2=1-q_1"

"1+q_2=-3+7q_2"

"q_2=2\/3" ",q_1=1\/3"

"V=1\/3+4\/3=5\/3"